OpenStudy (anonymous):

Solve: 3(x + 1) = −2(x − 1) − 4 Please help?

OpenStudy (anonymous):

use the factoring method. factor one side then the other.

OpenStudy (anonymous):

I tried to plug in all of the choices (1, -1, -5, -25) and I kept getting it wrong.

OpenStudy (anonymous):

\[3(x+1)=-2(x-1)-4\]\[3x+3=-2x+1-4\]\[3x+2x=1-4-3\]\[5x=-6\]\[x=\frac{ -6 }{ 5 }\]